# BlockingPy: approximate nearest neighbours for blocking of records for entity resolution

### v3

@@ Abstract
Entity resolution (probabilistic record linkage, deduplication) is a key step in scientific analysis and data science pipelines involving multiple data sources. The objective of entity resolution is to link records without common unique identifiers that refer to the same entity (e.g., person, company). However, without identifiers, researchers need to specify which records to compare in order to calculate matching probability and reduce computational complexity. One solution is to deterministically block records based on some common variables, such as names, dates of birth or sex or use phonetic algorithms. However, this approach assumes that these variables are free of errors and completely observed, which is often not the case. To address this challenge, we have developed a Python package, BlockingPy, which uses blocking using modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons. The package supports both CPU and GPU execution. In this paper, we present the design of the package, its functionalities and two case studies related to official statistics. The presented software will be useful for researchers interested in linking data from various sources.
@@
C2 Permanent link to code/repository used for this code version https://github.com/ncn-foreigners/BlockingPy
@@ 1 Motivation and significance
Entity resolution requires statistical and computational methodologies to accurately identify matching records across datasets without unique identifiers. This process underpins countless research endeavours in disciplines such as epidemiology, economics, official statistics and historical research where data integration is fundamental to scientific discovery [1, 2]. Traditional blocking techniques, while computationally efficient, exhibit limitations when confronted with various data quality issues such as missing values, transliteration, typographical errors, and inconsistent formatting. Existing deterministic blocking methods rely on user-defined exact-matching rules that assume key fields (e.g., name, date of birth) are error-free, potentially leading to missed links and false matches [3, 4, 5].
Scientific community has developed several tools to address record linkage, including R packages such as RecordLinkage [6], reclin2 [7], fastLink [8], or Python implementations like splink [9], recordlinkage [10], Dedupe [11] or FEBRL [12]. However, these solutions typically rely on user-defined deterministic rules that may miss certain groups of records or result in large blocks of records when data irregularities are present.
Some solutions regarding these issues were proposed in the literature, for instance, the use of the SOUNDEX algorithm and other phonetic algorithms to block records on character strings that sound similar phonetically [13]; blocking using the Locality Sensitive Hashing (LSH; [14]), an approximate nearest neighbour (ANN) algorithm to search for similar records (see klsh package in R) or usage of Bloom filters, useful in particular for privacy-preserving record linkage (see, [15]; for implementation see blocklib in Python).
The BlockingPy package uses state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data. BlockingPy implements an ANN-based blocking procedure ($n$-gram encoding or embeddings, neighbour search, graph components) behind a single pandas-style call, cutting custom rule-writing and setup time.
The structure of the paper is as follows. In Section 2, we briefly review the blocking procedures in existing Python packages. In Section 3, we provide details regarding the implementation and the API. In Section 4, we provide two examples of probabilistic record linkage and deduplication. The paper finishes with a brief review of the possible impact and conclusions.
@@ 2 Existing software covering blocking procedures
In this section, the focus is on Python, but readers are referred to the fastLink and reclin2 packages for more information on deterministic blocking procedures.
The splink package allows for deterministic blocking rules using the function block_on() which specifies either the variables (e.g. block_on("city")) or more advanced combinations (e.g. block_on("substr(first_name, 1,1)", "surname")), which then are translated to a SQL query for the DuckDB (via the BlockingRuleCreator class). More advanced users may be interested in creating their own rules, which is possible through the CustomRule class. Splink allows users to specify the list of blocking rules for the Linker class. Splink allows users to inspect the blocking via the
splink.blocking_analysis method.
Additionally, splink supports blocking using phonetic algorithms such as Soundex, Metaphone, or Double Metaphone either by applying them directly within the blocking rules (e.g., block_on("soundex(surname)")), or by preprocessing columns to store phonetic encodings separately. In online documentation we show that using splink with BlockingPy together results in significantly improvement in the performance metrics by capturing comparison pairs that would otherwise be missed.
@@
The only software that provides similar functionality is the blocklib package (cf. [16, 17]) and vicinity [18]. The first uses probability signature and LSH-based Lambda Fold Redundant as described in [16]. The second uses semhash [19], which allows deduplication of records using kNN-based embedding comparisons and offers exact-match blocking. In contrast, BlockingPy allows various inputs as well as ANN algorithms and graphs, focusing on providing accurate blocking for probabilistic record linkage and deduplication. In Appendix Appendix A. Numerical comparison with blocklib, we compare BlockingPy with blocklib in terms of computational efficiency, accuracy, and number of pairs from the blocking procedures. While blocklib is faster, it results in a significantly higher number of missed pairs and several orders of magnitude more pairs. We also provide comparison with the GPU version of the package.
@@ 3.2.1 Blocking records
The core functionality of BlockingPy is to provide an efficient and scalable approach for blocking (also known as indexing) records in both record linkage and deduplication pipelines. Scalability is achieved mainly due to the ANN libraries that we use in our implementations. Users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input. Alternatively, users can choose to represent the input data as vector embeddings.
@@
Users can fine-tune each algorithm according to their needs. Finally, BlockingPy groups records into blocks by identifying connected components in an undirected graph (using the igraph package [27]), generated by previous neighbourhood search results. By using the block() method with the desired parameters on the Blocker instance, users can run the above-mentioned process to obtain blocking results.
@@ 3.2.2 Evaluation of blocking result
Recall $\frac{TP}{TP+FN}$ Accuracy $\frac{TP+TN}{TP+TN+FP+FN}$
Precision $\frac{TP}{TP+FP}$ Specificity $\frac{TN}{TN+FP}$
@@ 4.1 Record linkage
Before blocking, we merge all fields of interest (without ID) into a single field (i.e., txt) that we will pass to the algorithm as shown in Listing 4. This step is performed for both datasets.
@@ 5 Impact
The proposed approach is used at Statistics Poland in the procedure for the linkage of records between administrative datasets as well as within one dataset. For instance, we have used this approach to deduplicate records of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022. We also used this technique to determine the number of residents of Ukraine under temporary protection in Poland as of March 31, 2023 [29]. The main motivation behind this was the problem of lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g. Олександр, should be Ołeksandr not Aleksandr). Before implementing this solution, Statistics Poland relied on user-defined rules for linking data without identifiers, which often resulted in imprecise population flow estimates due to missed matches or false positives.
@@ Acknowledgements
We thank the reviewers for their comments on this manuscript. This software is a part of Tymoteusz Strojny’s Bachelor’s thesis at Poznań University of Economics and Business, Poland. BlockingPy is based on the blocking package in R (see https://ncn-foreigners.github.io/blocking/).
@@
Appendix A. Numerical comparison with blocklib
We compare BlockingPy and blocklib in three synthetic datasets ($1,500$, $15,000$, and $150,000$ records) with duplicates $500$, $5,000$, and $50,000$, respectively. The results are averaged over 10 runs. The datasets were generated with geco3 and contain fictional personal information (given name, middle name, surname, municipality, date of birth, and nationality). CPU experiments used a 6-core Intel Core i5 with 16 GB RAM and Python 3.12. GPU runs used a NVIDIA RTX 3050 (4 GB VRAM) and Python 3.10. We report time (s), recall, reduction ratio (RR), and candidate pairs (millions). The comparison is also available in the documentation along with the code and data in the project repository. Preprocessing: for BlockingPy, we concatenate all fields of interest into a single text field; for blocklib, we tune algorithm parameters to approximately match the reduction ratio of the other methods to ease comparison.

### v4

@@ Abstract
Entity resolution (probabilistic record linkage, deduplication) is a key step in scientific analysis and data science pipelines involving multiple data sources. The objective of entity resolution is to link records without common unique identifiers that refer to the same entity (e.g., person, company). However, without identifiers, researchers need to specify which records to compare in order to calculate matching probabilities and reduce computational complexity. One solution is to deterministically block records based on some common variables, such as names, dates of birth, or sex, or by using phonetic algorithms. However, this approach assumes that these variables are free of errors and completely observed, which is often not the case. To address this challenge, we have developed a Python package, BlockingPy, which relies on schema-agnostic blocking using modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons. The package supports both CPU and GPU execution. In this paper, we present the design of the package, its functionalities and two case studies related to official statistics. The presented software will be useful for researchers interested in linking data from various sources.
@@
C2 Permanent link to code/repository used for this code version N/A
@@ 1 Motivation and significance
Entity resolution requires statistical and computational methodologies to accurately identify matching records across datasets without unique identifiers. This process underpins countless research endeavours in disciplines such as epidemiology, economics, official statistics and historical research where data integration is fundamental to scientific discovery [1, 30]. Traditional blocking techniques, while computationally efficient, exhibit limitations when confronted with various data quality issues such as missing values, transliteration, typographical errors, and inconsistent formatting. Existing deterministic blocking methods rely on user-defined exact-matching rules that assume key fields (e.g., name, date of birth) are error-free, potentially leading to missed links and false matches [4, 13, 14]. An alternative approach is schema-agnostic blocking [23, 31, 24], which concatenates all attribute values of a particular entity into a single textual representation before applying blocking (and can be combined with pre-trained language models).
The scientific community has developed several tools to address record linkage, including R packages such as RecordLinkage [25], reclin2 [32], fastLink [11], or Python implementations like splink [16], recordlinkage [9], Dedupe [12], pyJedAI [21], or FEBRL [3]. However, these solutions typically rely on user-defined deterministic rules that may miss certain groups of records or result in large blocks of records when data irregularities are present.
Some solutions to these issues were proposed in the literature, for instance, the use of the SOUNDEX algorithm and other phonetic algorithms to block records on character strings that sound similar phonetically [5]; blocking using the Locality Sensitive Hashing (LSH; [29]), an approximate nearest neighbour (ANN) algorithm to search for similar records (see, e.g., the klsh package in R) or the use of Bloom filters, useful in particular for privacy-preserving record linkage (see, [2]; for implementation see blocklib in Python).
The BlockingPy package relies on the schema-agnostic approach and uses state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data. BlockingPy implements an ANN-based blocking procedure ($n$-gram encoding or embeddings, neighbour search, graph components) behind a single pandas-style call, cutting custom rule-writing and setup time.
The structure of the paper is as follows. In Section 2, we briefly review the blocking procedures in existing Python packages. In Section 3, we provide details regarding the implementation and the API. In Section 4, we provide two examples of probabilistic record linkage and deduplication. The paper concludes with a brief review of the possible impact and conclusions.
@@ 2 Existing software covering blocking procedures
This section focuses on Python packages (splink, recordlinkage, and blocklib), but readers interested in deterministic blocking procedures are referred to the fastLink and reclin2 packages in R. For a comparison of Python packages, see Table 4 in the Appendix.
The splink package allows for deterministic blocking rules using the function block_on() which specifies either the variables (e.g., block_on("city")) or more advanced combinations (e.g., block_on("substr(first_name, 1,1)", "surname")), which then are translated to a SQL query for the DuckDB (via the BlockingRuleCreator class). More advanced users may be interested in creating their own rules, which is possible through the CustomRule class. Splink allows users to specify the list of blocking rules for the Linker class. Splink allows users to inspect the blocking via the splink.blocking_analysis method.
Additionally, splink supports blocking using phonetic algorithms such as Soundex, Metaphone, or Double Metaphone either by applying them directly within the blocking rules (e.g., block_on("soundex(surname)")), or by preprocessing columns to store phonetic encodings separately. In online documentation we show that using splink with BlockingPy together results in significant improvements in the performance metrics by capturing comparison pairs that would otherwise be missed.
@@
The only software that provides similar functionality is the blocklib package (cf. [34, 8]), vicinity [20], and pyJedAI. The first uses probability signature and LSH-based Lambda-Fold Redundant as described in [34]. The second uses semhash [33], which allows deduplication of records using kNN-based embedding comparisons and offers exact-match blocking. The third provides an end-to-end entity resolution pipeline that includes multiple key-based blocking methods and comparison cleaning (meta-blocking) strategies, with an optional embedding-based candidate generation component based on FAISS. In contrast, BlockingPy allows various inputs as well as ANN algorithms and graphs, focusing on providing accurate blocking for probabilistic record linkage and deduplication. In B, we compare BlockingPy with blocklib in terms of computational efficiency, accuracy, and number of pairs from the blocking procedures. While blocklib is faster, it results in a significantly higher number of missed pairs and several orders of magnitude more pairs. We also provide a comparison with the GPU version of the package. In C we compare PyJedAI with BlockingPy on the Abt-Buy benchmark [15].
@@ 3.2.1 Blocking records
The core functionality of BlockingPy is to provide an efficient and scalable approach for blocking (also known as indexing) records in both record linkage and deduplication pipelines. Scalability is achieved mainly due to the ANN libraries that we use in our implementations. Users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input. Alternatively, users can represent the input data as vector embeddings. For models supported by the model2vec library, users can simply pass the model identifier as a string (e.g., "minishlab/potion-base-8m"), and the package computes the embeddings internally. For any other embedding model, users can precompute the embeddings separately and pass them to the package as NumPy arrays.
@@
Users can fine-tune each algorithm according to their needs. Finally, BlockingPy groups records into blocks by identifying connected components in an undirected graph (using the igraph package [6]), generated by previous neighbourhood search results. In this graph, vertices represent records and edges connect pairs of records identified as neighbours. This step is necessary to handle transitive relationships (e.g., record A matches record B and record B matches record C). By using the block() method with the desired parameters on the Blocker instance, users can run the above-mentioned process to obtain blocking results.
@@ 3.2.2 Evaluation of blocking result
Recall (Pairs Completeness) $\frac{TP}{TP+FN}$ Accuracy $\frac{TP+TN}{TP+TN+FP+FN}$
Precision (Pairs Quality) $\frac{TP}{TP+FP}$ Specificity $\frac{TN}{TN+FP}$
@@ 4.1 Record linkage
Before blocking, we merge all fields of interest (without ID) into a single field (i.e., txt) that we will pass to the algorithm as shown in Listing 4 (i.e, we use schema-agnostic blocking). This step is performed for both datasets.
@@ 5 Impact
The proposed approach is used at Statistics Poland in the procedure for the linkage of records between administrative datasets as well as within one dataset. For instance, we have used this approach to deduplicate records of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022. We also used this technique to determine the number of residents of Ukraine under temporary protection in Poland as of March 31, 2023 [28]. The main motivation behind this was the problem of the lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g. Олександр, should be Ołeksandr not Aleksandr). Before implementing this solution, Statistics Poland relied on user-defined rules for linking data without identifiers, which often resulted in imprecise population flow estimates due to missed matches or false positives.
@@ Acknowledgements
We thank the reviewers for their comments on this manuscript. This software was a part of Tymoteusz Strojny’s Bachelor’s thesis at Poznań University of Economics and Business, Poland. BlockingPy is based on the blocking package in R (see https://ncn-foreigners.github.io/blocking/).
@@
Appendix A Comparison of blocking methods implemented in splink, blocklib, recordLinkage, pyJedi and BlockingPy
Table 4: Comparison of blocking methods implemented in selected software.
Package Blocking methods
splink Rule-based / deterministic blocking (SQL-style rules), phonetic blocking (e.g., Soundex, Metaphone).
blocklib Probabilistic signature blocking (p-sig), LSH-based $\lambda$-fold redundant blocking (Bloom filter + LSH).
recordlinkage Exact blocking, Sorted Neighbourhood, full indexing (all pairs), random indexing (sampled pairs).
pyJedAI Token Blocking, Sorted Neighborhood, Extended Sorted Neighborhood, Q-Grams Blocking, Extended Q-Grams Blocking, Suffix Arrays Blocking, Extended Suffix Arrays Blocking, Embeddings-based NN blocking.
BlockingPy ANN-based blocking via FAISS, HNSW (faiss / hnswlib / voyager), NN-Descent (pynndescent), Annoy, plus GPU indices (e.g., IVF, IVFPQ, CAGRA); blocks from connected components of the NN graph.
Appendix B Numerical comparison with blocklib
We compare BlockingPy and blocklib in three synthetic datasets ($1,500$, $15,000$, and $150,000$ records) with duplicates $500$, $5,000$, and $50,000$, respectively. The results are averaged over 10 runs. The datasets were generated with geco3 and contain fictional personal information (given name, middle name, surname, municipality, date of birth, and nationality). CPU experiments used a 6-core Intel Core i5 with 16 GB RAM and Python 3.12. GPU runs used a NVIDIA RTX 3050 (4 GB VRAM) and Python 3.10. We report time (s), recall (Pairs Completeness), reduction ratio (RR), and candidate pairs (millions). The comparison is also available in the documentation along with the code and data in the project repository. Preprocessing: for BlockingPy, we concatenate all fields of interest into a single text field; for blocklib, we tune algorithm parameters to approximately match the reduction ratio of the other methods to ease comparison.
@@ Appendix B Numerical comparison with blocklib
Appendix C Comparison with PyJedAI
We compare BlockingPy and PyJedAI on the Abt-Buy clean-clean entity resolution benchmark, which consists of two product catalogs from the online retailers Abt.com and Buy.com. We use the Abt-Buy dataset as provided in the pyJedAI repository, which contains 1,076 records in each source and a gold-standard mapping between the two sources. The hardware setup is the same as in B; the experiments were run with Python 3.11 (with PyJedAI using NumPy 2.4.1 and BlockingPy using NumPy 1.26.4). We report runtime (s), recall, precision, $F_{1}$, and the number of candidate comparison pairs.
For BlockingPy, we use the HNSW backend with the minishlab/potion-base-32M embedding model; all remaining parameters are left at their defaults. For PyJedAI, we use the sminilm vectorizer with $k=5$ nearest neighbors and Euclidean distance. The data (and an example workflow) are available in the pyJedAI documentation/repository. The BlockingPy workflow used for this comparison is likewise available in the BlockingPy repository.
Table 7: Comparison on the Abt-Buy benchmark
Algorithm Time [s] Recall Precision $F_{1}$ Pairs PyJedAI 52.6 0.9377 0.1875 0.3126 5380 BlockingPy 4.4 0.8234 0.8234 0.8234 1076
Note: The identical recall, precision, and $F_{1}$ values reported here are expected in this setting and are not a reporting error: in BlockingPy we use $k=1$ (one candidate per record) and, for this experiment, the graph-based step did not find additional links, so evaluation is effectively performed on the same one-to-one pair set; under these conditions these three metrics coincide.